LINE MIXING EFFECTS IN THE ν_{2}+ν_{3} BAND OF METHANE.

ADRIANA PREDOI-CROSS, ANIL V. UNNIKRISHNAN, HENRY HEUNG, University of Lethbridge, Department of Physics, 4401 University Drive, Lethbridge, Alberta, TIK3M4, Canada; \underline{LINDA R. BROWN}, Jet Propulsion Laboratory, California Institute of Technology, Pasadena, California 91109; D. CHRIS BENNER and V. MALATHY DEVI, Department of Physics, The College of William and Mary, Williamsburg, VA 23187-8795.

This study provides the first experimental measurements of line mixing via the off diagonal relaxation matrix element formalism in air-broadened methane spectra for any vibrational band and the first off diagonal relaxation matrix elements associated with line mixing for pure methane in the ν_{2}+ν_{3} band of ^{12}CH_{4}. A speed-dependent Voigt profile with line mixing is used with a multispectrum nonlinear least squares curve fitting technique.^{a} The off diagonal relaxation matrix element coefficients of eighteen pairs of ν_{2}+ν_{3} transitions between 4410 and 4629 cm^{−1} have been determined. The measured self-line mixing coefficients vary from 0.0019 to 0.0390 cm^{−1}atm^{−1} at 296 K, and for air line mixing coefficients vary between 0.0005 and 0.0205 cm^{−1}atm^{−1} at 296 K. The spectral data used in the analysis were recorded at a resolution of 0.01-cm^{−1} using the McMath-Pierce Fourier transform spectrometer located at the National Solar Observatory on Kitt Peak, Arizona.

aD. Chris Benner et al. JQSRT \underline{53}, 705-721, 1995.